Grade 12th passAlgebra

Write a Pythagorean triplet whose one member is. . 16

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

2m=16⇒m=16/2=8

m2–1=82–1=64–1=63

m2+1=82+1=64+1=65
∴ (16, 63, 65) is aPythagorean triplet.

Thanks